A discrete random variable x can assume five possible values: 2, 3, 5, 8, 10. It’s probability distribution is shown below: Find the probability that the random variable x value  greater than 5

Probability Distribution:

x 2 3 5 8 10

P (x) 0.100.2.00.300.300.10
Answer #1

Solution:-

Given data:-

P(X=2)=0.10

P(X=3)=0.20

P(X=5)=0.30

P(X=8)=0.30

P(X=10)=0.10

P(X>5)=P( X = 8) + P( X = 10)

where,P(X=8)=0.30 and P(X=10)=0.10

P(X>5)=0.4